# AD-L-JEPA: Self-Supervised Spatial World Models with Joint Embedding Predictive Architecture for Autonomous Driving with LiDAR Data

### v1

@@ Abstract
As opposed to human drivers, current autonomous driving systems still require vast amounts of labeled data to train. Recently, world models have been proposed to simultaneously enhance autonomous driving capabilities by improving the way these systems understand complex real-world environments and reduce their data demands via self-supervised pre-training. In this paper, we present AD-L-JEPA (aka Autonomous Driving with LiDAR data via a Joint Embedding Predictive Architecture), a novel self-supervised pre-training framework for autonomous driving with LiDAR data that, as opposed to existing methods, is neither generative nor contrastive. Our method learns spatial world models with a joint embedding predictive architecture. Instead of explicitly generating masked unknown regions, our self-supervised world models predict Bird’s Eye View (BEV) embeddings to represent the diverse nature of autonomous driving scenes. Our approach furthermore eliminates the need to manually create positive and negative pairs, as is the case in contrastive learning. AD-L-JEPA leads to simpler implementation and enhanced learned representations. We qualitatively and quantitatively demonstrate high-quality of embeddings learned with AD-L-JEPA. We furthermore evaluate the accuracy and label efficiency of AD-L-JEPA on popular downstream tasks such as LiDAR 3D object detection and associated transfer learning. Notably, our method achieves up to a 5×5$\times$5 × speed-up in pre-training time compared with the state-of-the-art (SOTA) approaches and consistently delivers better performance when fine-tuned on different amounts of labeled data compared to SOTA. More impressively, the experiments show that even when the first layers of the encoders of AD-L-JEPA are randomly initialized while the remaining layers are initialized as pre-trained, our method still yields positive transfers, whereas the competitor method suffer from the negative transfers in this case. Systematic experimental results demonstrate that our method learns a robust spatial world model in the BEV embedding space, resulting in faster pre-training times, improved generalization, and enhanced label efficiency. Our experimental evaluation demonstrates that AD-L-JEPA is a plausible approach for self-supervised pre-training in autonomous driving applications and is the best available approach outperforming SOTA, including most recently proposed Occupancy-MAE [1] and ALSO [2]. The source code of AD-L-JEPA is available at https://github.com/HaoranZhuExplorer/AD-L-JEPA-Release.
I INTRODUCTION
Figure 1: AD-L-JEPA consistently improves label efficiency when transferring from the Waymo (subsampled $20\%$ of the data set is used) to the KITTI3D data set in the LiDAR object detection task (we use SECOND [3] detection model), and achieves a 5×5$\times$5 × speed-up in pre-training time.
@@ I INTRODUCTION
LiDAR data is one of the most common data modalities in autonomous driving. Existing autonomous driving systems need to collect vast amounts of raw LiDAR data across various driving scenarios and then annotate them with labels, which is expensive, time-consuming, and limits the scalability. And in particular, label annotation constitutes a significant bottleneck here since the collection of unlabeled data alone is much easier. Recently, researchers have proposed to have a common backbone for different learning tasks that is pre-trained in a self-supervised manner without any labels and then only fine-tune the whole network with labeled data [1, 2] to adapt to specific downstream tasks. This approach requires less labeled data and is therefore more practical.
A key to further enhancing autonomous driving capabilities may therefore lie in learning self-supervised world models [4] as a novel pre-training paradigm. Furthermore, learning world models in a self-supervised manner enables the network to predict unknown states of the environment, which helps to manage uncertainty in the autonomous driving systems.
What is our contribution and what makes us different from SOTA? In this paper, we propose a novel self-supervised pre-training method AD-L-JEPA for LiDAR data. To the best of our knowledge, this is the first JEPA-based pre-training framework for autonomous driving. Our method learns a self-supervised spatial world model in Bird’s Eye View (BEV) space and predicts embeddings for spatially unknown regions. Since it predicts BEV embeddings, instead of explicitly reconstructing the unknown parts of the data, it improves the quality of the data modeling (embeddings contain the most relevant information, including relevant details, for example a car hidden behind the truck may not show in reconstructed data but still can be well captured in BEV embeddings) and allows to better adapt to the data uncertainty, which leads to better generalization. It is also extremely efficient and significantly reduces the pre-training time. Furthermore, AD-L-JEPA learns high-quality embeddings, which also results in superior label efficiency on downstream tasks. Motivating illustration demonstrating the advantages of AD-L-JEPA is captured in Figure 1. The implementation of AD-L-JEPA is straightforward and omits the necessity to create human-crafted positive/negative pairs, as is the case in contrastive learning. Finally, AD-L-JEPA adds to the existing family of self-supervised pre-training methods for autonomous driving applications a new approach that compares favorably to the existing SOTA, including most recently proposed Occupancy-MAE [1] and ALSO [2]. The source code of AD-L-JEPA is released.
II BACKGROUND AND RELATED WORKS
II-A Self-supervised learning
@@
Self-supervised learning aims at learning useful data representations from the unlabeled data. We focus here on self-supervised learning appraoches that are dedicated to autonomous driving with LiDAR point cloud data. Among the existing methods, contrastive techniques learn self-supervised representations by maximizing the feature similarity of manually created positive pairs (matching points) from different levels, such as point level [5], depth map level [6], region clusters level [7], spatial or temporal segments [8, 9, 10, 11, 12, 13, 14, 15], or Bird’s Eye View (BEV) features over time [16], while minimizing the similarity of negative pairs (non-matched points), if any. Another family of approaches are the generative-based method, where the network is trained to reconstruct masked point clouds [1, 17, 18, 14] or scene surfaces [2, 19].
@@
Besides contrastive and generative methods, another novel self-supervised learning paradigm offers joint-embedding predictive architecture (JEPA). It has recently been applied to image and video data modalities [4, 20, 21] but has not yet been explored for autonomous driving scenarios. JEPA learns meaningful representations by predicting the embeddings of the unknown parts of data extracted by a target encoder given the known parts embeddings extracted by a context encoder. Unlike contrastive learning, it does not require complicated pre-processing to create positive pairs and does not suffer from the curse of dimensionality with negative pairs (these are the commonly-known drawbacks of contrastive learning [4]). Furthermore, it effectively captures the high uncertainty of the environment without explicitly generating reconstructions of unknown regions that is typically done by generative methods. Our proposed AD-L-JEPA is the first JEPA-based pre-training method dedicated for autonomous driving.
@@
Representation collapse is a common phenomenon, when the network fails to learn meaningful representations. It can manifest in two ways: complete collapse, where all representations reduce to a constant vector and thus become useless, or dimensional collapse, where the dimensions of learned representations are highly correlated with each other and contain redundant information. Generative-based methods naturally avoid complete collapse when the dimension of the information bottleneck in the encoder is smaller than the dimension of the input, as generating unknown regions requires learning meaningful representations. Contrastive-based methods avoid complete collapse by creating contrastive pairs to learn distinctive representations. For JEPA-based methods, regularization techniques and moving average updates of the target encoder have been proven useful [22, 20, 21]. There are also studies that quantitatively measure the degree of representation collapse by examining the singular value spectrum of the learned representations [23]. In our paper, we employ both regularization techniques and moving average updates of the target encoder to prevent collapse.
II-B World models for autonomous driving
World Models [24, 4, 25] aim at modeling the possible spatially unknown states or future states of the environment to help the agent navigate in it. World models help understanding the data and model uncertainties in the data. Instead of explicitly predicting data in pixel space, the method based on the joint-embedding predictive architecture [20] proposes predicting embeddings of masked or future image patches in a self-supervised manner [21] to capture multiple potential states. Also, given a context video or text, world models allow to generate possible future videos. In autonomous driving scenarios, world models have been deployed as a data engine [26], to mitigate covariance shift [27], to improve end-to-end driving [28, 29], and for pre-training [30].
II-C LiDAR-based 3D object detection
@@
We use 3D Object Detection as the downstream task to evaluate the effectiveness of our pre-training method for autonomous driving with LiDAR data. For most LiDAR-based 3D object detection downstream algorithms, a sparse 3D convolution encoder first extracts 3D voxel representations from a given point cloud scene. It then reshapes the image over the height dimension and applies 2D dense convolutions to generate BEV representations. Finally, either a single-stage [3] or a two-stage detection head [31, 32, 33] is attached for the LiDAR-based 3D object detection task. Following common settings in the literature, we focus on conducting self-supervised pre-training and evaluate the pre-trained feature quality also using the sparse 3D convolution encoder framework.
III METHOD
Figure 2: The overview of AD-L-JEPA architecture.
@@ III METHOD
The architecture of AD-L-JEPA is shown in Figure 2. The overarching intuition behind our framework is as follows: for the visible parts of the point cloud scene, the network is trained in a self-supervised manner to predict how the invisible parts should appear in the embedding space. This enables the learning of geometrically and semantically reasonable world models, as well as adapting to the high uncertainty nature of the autonomous driving scenes by avoiding the explicit reconstruction of the invisible parts of the data.
In the following subsections we describe our design in detail. First, we mask the input point cloud with our proposed modified BEV-guided masking (Section III-A). Next, unmasked points are sent to the context encoder, and masked points are sent to the target encoder (Section III-B). We then replace the empty regions in both context and target BEV embeddings with a learnable empty token, and the masked regions in the context BEV embedding with a learnable masked token (Section III-C). Later, a lightweight predictor predicts target BEV embeddings from context BEV embeddings (Section III-D). Finally, we train the context encoder and predictor using cosine similarity-based embedding prediction loss and variance regularization loss, while we update the target encoder through a moving average strategy. This ensures the learning of high-quality, non-collapsed representations (Section III-E).
III-A Modified BEV-guided masking
@@
To learn effective spatial world models in a self-supervised manner, masking is used to create invisible and visible regions. The network is then trained to predict embeddings of the invisible regions based on the visible ones. We have two design recipes for masking in autonomous driving scenarios: (1) masks are first created in the BEV embedding space and recursively upsampled to the input point cloud to identify points to be masked; (2) both empty and non-empty areas should be included in the visible and invisible regions created by the masks. These two criteria can be achieved by modifying the BEV-guided masking originally proposed in [17] (see comparison in Figure 3).
@@
Figure 3: Comparison of original BEV-guided masking [17] with our modified version that creates masks in both empty and non-empty regions in non-overlapping BEV grids.
@@
The original BEV-guided masking involves projecting points onto non-overlapping BEV grids, creating masks only in non-empty BEV grids, and reconstructing only in the masked non-empty regions. This assumes that the network already knows which BEV grid is empty, which stands in contrast with our goal to learn a spatial world model to predict all invisible regions. As opposed to this approach, we create masks for both empty and non-empty BEV grids, resulting in visible and invisible parts that contain both empty and non-empty regions. Consequently, the network is trained to predict the embeddings of all invisible parts, including those that contain empty regions, which enhances the encoder’s representation power and helps to learn good quality spatial world models. In our experiments, masks are applied to $50\%$ of non-empty BEV grids and $50\%$ of empty BEV grids. The input points sent to the encoder remain the same as in the original BEV-guided masking. However, the encoder’s output BEV embeddings will later be replaced by mask tokens for both masked empty and non-empty grids, whereas in [17], only masked non-empty grids were replaced by mask tokens. We denote the unmasked point cloud as $\boldsymbol{x}_{c}=\{\boldsymbol{x}^{1}_{c},\ldots,\boldsymbol{x}^{N}_{c}\}$ and the masked point cloud as $\boldsymbol{x}_{t}=\{\boldsymbol{x}^{1}_{t},\ldots,\boldsymbol{x}^{N}_{t}\}$ in the multi-batch setting with a batch size of $N$. $\boldsymbol{x}_{c}$ and $\boldsymbol{x}_{t}$ are then sent to the context encoder and target encoder, respectively.
III-B Context encoder & target encoder
@@
III-C Learnable empty token and mask token
@@ III-E Training
where $P_{n}$ is the subset containing all indices of unmasked empty BEV grids, and $Q_{n}$ is the subset containing all indices of unmasked non-empty BEV grids in the $n$-th input in a multi-batch setting. We also denote $K_{n}$ as the subset containing all indices of masked non-empty BEV grids in the $n$-th input.
In order to avoid representation collapse, we apply an additional variance regularization loss proposed by [22]:
@@
where the function $v(\cdot)$, which uses an arbitrary input 2D embedding matrix $Y\in\mathbb{R}^{M\times C}$ recording an arbitrary number $M$ of embeddings with a fixed embedding dimension $C$, is defined in [22]:
@@
The parameters of the context encoder, the predictor, as well as the learnable mask token and the learnable empty token, are all updated by gradient descent using Equation 4. Meanwhile, the parameters of the target encoder are updated through a moving average of the context encoder’s parameters, $\phi\leftarrow\eta\phi+(1-\eta)\theta$, to further avoid representation collapse, as proposed in [20, 21].
IV EXPERIMENTS
To evaluate our pre-training method relying on self-supervised spatial world models, we use a LiDAR 3D object detection task as a downstream fine-tuning task. We utilize the OpenPCDet framework [34] to pre-process the data sets and conduct the experiments.
IV-A Data sets
We use KITTI3D [35] and Waymo [36] data sets to conduct self-supervised pre-training and downstream fine-tuning. The KITTI3D data set contains a total of $3712$ training frames and $3769$ validation frames. We report the $3$D object detection Average Precision (AP) metric, calculated at $40$ recall positions ($R_{40}$) and $11$ recall positions ($R_{11}$), for the moderate level of difficulty  [^1] Moderate level of difficulty is defined in [35] and LEVEL 2 difficulty, mentioned later in the text, is defined in [36]. across different classes on KITTI3D data set. The Waymo data set contains a total of $158081$ training frames and $39987$ validation frames. Following the popular setting of OpenPCDet, we subsample $20\%$ of the training frames with a sampling interval of $5$, resulting in $31617$ training frames. We report the $3$D object detection Average Precision (AP) metric and Average Precision with Heading (APH) metric for LEVEL 2 difficulty across different classes on Waymo data set.
IV-B Network settings
We use our proposed AD-L-JEPA to pre-train the popular VoxelBackBone8x encoder, which contains $12$ sparse convolution layers. For both the KITTI3D and $20\%$ subsampled Waymo data sets, we pre-train the network with unlabeled training data only for $30$ epochs using the Adam optimizer and the default one-cycle learning rate scheduler in OpenPCDet, with a learning rate of $0.0003$ and a weight decay of $0.01$. We set the initial $\eta$ to 0.996 and increase it linearly to $1$, following [21], for the moving average update. We set masking ratio = $0.5$. We set $\alpha_{0}=0.25$ and $\alpha_{1}=0.75$ for empty grids and non-empty, respectively. We set $\beta_{1}=1$ and $\beta_{2}=1$. The threshold for variance regularization, $\gamma$, is set to $\frac{1}{\sqrt{D}}=\frac{1}{16}$ for $L_{2}$-normalized embeddings. For the KITTI3D data set, the shape for BEV embedding is $H=200$, $W=176$, $D=2$, $C=128$; we set $\lambda_{\text{jepa}}=1$ and $\lambda_{\text{reg}}=1$ to balance the two loss functions. For the Waymo data set, the shape for BEV embedding is $H=188$, $W=188$, $D=2$, $C=128$; we set $\lambda_{\text{jepa}}=1$ and $\lambda_{\text{reg}}=10$ to balance the two loss functions. We use $4$ $1080$ Ti GPUs for KITTI3D and a single $A100$ GPU for Waymo during pre-training, with a total batch size of $16$.
@@ IV-B Network settings
During fine-tuning, we initialize the encoder’s weights with the pre-trained context encoder’s weights and then fine-tune using detection models: SECOND [3], PVRCNN [33], and Centerpoint [32] (we use encoders without residual connections) for $80$ epochs, with all default hyperparameter settings in the OpenPCDet framework. We compare models trained from scratch and the most recent SOTA self-supervised pre-training methods, Occupancy-MAE [1] and ALSO [2]. Specifically, we reproduce Occupancy-MAE using their default settings to obtain the pre-trained weights and fine-tune the downstream object detector. For ALSO, we directly fine-tune the released pre-trained encoder weights for the downstream object detector. The comparison methods use the same fine-tuning settings as ours. For all fine-tuning experiments, we fine-tune the model through three independent runs and report the best performance, in the same manner as [16] reports their results. The reproduced results may vary from the original papers. We use $4$ $1080$ Ti GPUs for KITTI3D and $4$ RTX $4090$ GPU for Waymo during fine-tuning, with a total batch size of $16$.
IV-C Downstream fine-tuning performance
@@
TABLE I: 3D detection results on the KITTI3D validation set with moderate difficulty, reported with the AP (%) metric. The models are pre-trained with and fine-tuned on the KITTI3D data set.
masking ratio 0.25 0.5 0.75
mAP - $R_{40}$ 66.21 67.92 67.24
mAP - $R_{11}$ 65.73 67.10 66.76
TABLE II: Different masking ratio and the corresponding mAP (%) on KITTI3D data set with SECOND detection model.
@@
Centerpoint (30 epoch) - AP metric
No pre-training 63.28 63.95 66.78 64.67
Occupancy-MAE [1] 63.20 64.20 67.20 64.87 +0.20
AD-L-JEPA (ours) 63.18 64.35 67.68 65.07 +0.40
Centerpoint (30 epoch) - APH metric

### v2

# Self-Supervised Representation Learning with Joint Embedding Predictive Architecture for Automotive LiDAR Object Detection

@@ Abstract
Recently, self‑supervised representation learning relying on vast amounts of unlabeled data has been explored as a pre‑training method for autonomous driving. However, directly applying popular contrastive or generative methods to this problem is insufficient and may even lead to negative transfer. In this paper, we present AD‑L‑JEPA, a novel self‑supervised pre‑training framework with a joint embedding predictive architecture (JEPA) for automotive LiDAR object detection. Unlike existing methods, AD‑L‑JEPA is neither generative nor contrastive. Instead of explicitly generating masked regions, our method predicts Bird’s‑Eye‑View embeddings to capture the diverse nature of driving scenes. Furthermore, our approach eliminates the need to manually form contrastive pairs by employing explicit variance regularization to avoid representation collapse. Experimental results demonstrate consistent improvements on the LiDAR 3D object detection downstream task across the KITTI3D, Waymo, and ONCE datasets, while reducing GPU hours by $1.9\times$$2.7\times$ and GPU memory by $2.8\times$$4\times$ compared with the state-of-the-art method Occupancy-MAE. Notably, on the largest ONCE dataset, pre‑training on 100K frames yields a 1.61 mAP gain, better than in case of all the other methods pre‑trained on either 100K or 500K frames, and pre‑training on 500K frames yields a 2.98 mAP gain, better than in case of all the other methods pre‑trained on either 500K or 1M frames. AD‑L‑JEPA constitutes the first JEPA‑based pre‑training method for autonomous driving. It offers better quality, faster, and more GPU‑memory‑efficient self‑supervised representation learning. The source code of AD-L-JEPA is ready to be released.
Figure 1: AD‑L‑JEPA predicts directly in embedding space instead of explicitly reconstructing masked point clouds, as in driving scenes these can correspond to multiple plausible point clouds sharing the same semantics (e.g., “car rear”) and can be encoded into the same embedding. It significantly boosts downstream performance while reducing pre-training time and GPU memory usage.
Introduction
@@
Unlike human drivers, current autonomous driving systems still require large amounts of labeled data for training. This supervised‑only paradigm is expensive due to labeling costs and limits the scalability of these systems. Recently, researchers have proposed a learning paradigm that relies on having a common backbone for different learning tasks that is pre-trained with self-supervised learning (SSL) across camera, LiDAR, and radar modalities (Yang et al. 2024; Min et al. 2023; Zhu et al. 2024) without any labels and then fine-tuning the whole network with labeled data to adapt to specific downstream tasks. This approach requires less labeled data and improves generalization capability.
In SSL, the two most popular learning paradigms are contrastive methods (He et al. 2020; Chen et al. 2020) and generative methods (He et al. 2022). However, directly applying these methods for pre‑training in autonomous driving is challenging and can even hurt downstream performance (Mao et al. 2021). This stems from both the difficulty of defining meaningful contrastive pairs via data augmentation in driving scenarios that contain multiple objects, and the fact that explicit scene generation is time‑consuming and insufficient to capture semantic representations of diverse driving scenarios.
In this paper, we present AD-L-JEPA (aka Autonomous Driving with LiDAR data via a Joint Embedding Predictive Architecture (LeCun 2022)), a novel self-supervised pre-training framework for automotive LiDAR object detection that, as opposed to existing methods, is neither generative nor contrastive. Our method learns self‑supervised representations in Bird’s Eye View (BEV) space and predicts embeddings for spatially masked regions. It omits the need to create human‑crafted positive/negative pairs, as required by contrastive learning. Furthermore, rather than explicitly reconstructing unknown parts of the data as generative methods do, it predicts BEV embeddings instead.
AD‑L‑JEPA improves the quality of data representations, capturing high‑level semantics and better adapting to uncertainty across multiple plausible driving scenarios (toy illustration capturing that AD-L-JEPA can nicely handle data uncertainty is shown in Figure 1, where different car rears for masked regions are plausible given the visible car front), leading to better generalization. It is also extremely efficient and significantly reduces the pre-training time and GPU memory usage. Extensive experiments show that AD-L-JEPA learns high-quality embeddings, which result in a consistent boost in fine-tuning and transfer learning and offer superior label efficiency for downstream tasks. A motivating illustration of AD‑L‑JEPA’s advantages is shown in Figure 1. Finally, AD‑L‑JEPA introduces a new approach to the existing family of self‑supervised pre‑training methods for autonomous driving, which compares favorably to the state‑of‑the‑art methods: Occupancy‑MAE (Min et al. 2023) and ALSO (Boulch et al. 2023). To the best of our knowledge, this is the first JEPA‑based pre‑training framework for autonomous driving.
Related Work
Self-Supervised Learning
@@
Self-supervised learning (Balestriero et al. 2023; He et al. 2020; Chen et al. 2020; Grill et al. 2020; Chen and He 2021; Caron et al. 2021; Zbontar et al. 2021; Bardes, Ponce, and LeCun 2022; He et al. 2022; Assran et al. 2023; Bardes et al. 2024; Assran et al. 2025) aims at learning useful data representations from the unlabeled data. We focus here on SSL appraoches that are dedicated to autonomous driving with LiDAR point cloud data. Among the existing methods, contrastive techniques learn self-supervised representations by maximizing the feature similarity of manually created positive pairs (matching points) from different levels, such as point level (Xie et al. 2020), depth map level (Zhang et al. 2021), region clusters level (Yin et al. 2022), spatial or temporal segments (Liang et al. 2021; Huang et al. 2021; Nunes et al. 2022; Wu et al. 2023; Nunes et al. 2023; Yuan et al. 2024; Wei et al. 2025; Hegde et al. 2025), or Bird’s Eye View (BEV) features over time (Sautier et al. 2024), while minimizing the similarity of negative pairs (non-matched points), if any. Another family of approaches are the generative-based method, where the network is trained to reconstruct masked point clouds (Min et al. 2023; Lin et al. 2024; Xu et al. 2023; Wei et al. 2025; Abdelsamad et al. 2025) or scene surfaces (Boulch et al. 2023; Agro et al. 2024).
@@
Besides contrastive and generative methods, another novel self-supervised learning paradigm offers joint-embedding predictive architecture (JEPA). It has recently been applied to image and video data modalities (LeCun 2022; Assran et al. 2023; Bardes et al. 2024; Assran et al. 2025) but has not yet been explored for autonomous driving scenarios. JEPA learns meaningful representations by predicting the embeddings of the unknown parts of data extracted by a target encoder given the known parts embeddings extracted by a context encoder. Unlike contrastive learning, it does not require complicated pre-processing to create positive pairs and does not suffer from the curse of dimensionality with negative pairs (LeCun 2022). Furthermore, it effectively captures the high uncertainty of the environment without explicitly generating reconstructions of unknown regions that is typically done by generative methods.
@@
Representation collapse is a common phenomenon, when the network fails to learn meaningful representations. It can manifest in two ways: complete collapse, where all representations reduce to a constant vector and thus become useless, or dimensional collapse, where the dimensions of learned representations are highly correlated with each other and contain redundant information. Introducing contrastive learning, or applying regularization techniques (Bardes, Ponce, and LeCun 2022) and moving average updates of the target encoder have been proven useful (Grill et al. 2020). In our paper, we employ both regularization techniques and moving average updates of the target encoder to prevent collapse.
LiDAR-Based 3D Object Detection
@@
We use 3D Object Detection as the downstream task to evaluate the effectiveness of our pre-training method for autonomous driving with LiDAR data. For most LiDAR-based 3D object detection downstream algorithms, a sparse 3D convolution encoder (Graham and Van der Maaten 2017) first extracts 3D voxel representations from a given point cloud scene. It then reshapes the image over the height dimension and applies 2D dense convolutions to generate BEV representations. Finally, either a single-stage (Yan, Mao, and Li 2018) or a two-stage detection head (Shi, Wang, and Li 2019; Yin, Zhou, and Krahenbuhl 2021; Shi et al. 2020; Deng et al. 2021) is attached for the LiDAR-based 3D object detection task. Following common settings in the literature, we focus on conducting self-supervised pre-training and evaluate the pre-trained feature quality also using the sparse 3D convolution encoder framework.
Figure 2: Overview of the AD-L-JEPA architecture: We introduce modified BEV-guided masking to mask the input point cloud in both empty and non-empty regions. The network predicts BEV embeddings at masked regions, leveraging variance regularization at non-empty regions following the output of the context encoder and the lightweight spatial predictor. It also employs a moving average update of the target encoder to learn diverse, high-level semantic representations.
Method
@@
The architecture of AD-L-JEPA is shown in Figure 2. The overarching intuition behind our framework is as follows: for the visible parts of the point cloud scene, the network is trained in a self-supervised manner to predict how the invisible parts should appear in the embedding space. This enables the learning of geometrically and semantically reasonable representations, as well as adapting to the high uncertainty nature of the autonomous driving scenes by avoiding the explicit reconstruction of the invisible parts of the data. In the following subsections we describe our design in detail.
Modified BEV-Guided Masking
@@
To learn effective representations in a self-supervised manner, masking is used to create invisible and visible regions. The network is then trained to predict embeddings of the invisible regions based on the visible ones. We have two design recipes for masking in autonomous driving scenarios: (1) masks are first created in the BEV embedding space and recursively upsampled to the input point cloud to identify points to be masked; (2) both empty and non-empty areas should be included in the visible and invisible regions created by the masks. These two criteria can be achieved by modifying the BEV-guided masking originally proposed in (Lin et al. 2024) (see comparison in Figure 3).
@@
Figure 3: Comparison of original BEV-guided masking (Lin et al. 2024) with our modified version that creates masks in both empty and non-empty regions in non-overlapping BEV grids.
@@
The original BEV-guided masking involves projecting points onto non-overlapping BEV grids, creating masks only in non-empty BEV grids, and reconstructing only in the masked non-empty regions. This assumes that the network already knows which BEV grid is empty, which stands in contrast with our goal to learn representations to predict all invisible regions. As opposed to this approach, we create masks for both empty and non-empty BEV grids, resulting in visible and invisible parts that contain both empty and non-empty regions. Consequently, the network is trained to predict the embeddings of all invisible parts, including those that contain empty regions, which enhances the encoder’s representation power and helps to learn good quality representations. In our experiments, masks are applied to $50\%$ of non-empty BEV grids and $50\%$ of empty BEV grids. The input points sent to the encoder remain the same as in the original BEV-guided masking. However, the encoder’s output BEV embeddings will later be replaced by mask tokens for both masked empty and non-empty grids, whereas in (Lin et al. 2024), only masked non-empty grids were replaced by mask tokens. We denote the unmasked point cloud as $\boldsymbol{x}_{c}=\{\boldsymbol{x}^{1}_{c},\ldots,\boldsymbol{x}^{N}_{c}\}$ and the masked point cloud as $\boldsymbol{x}_{t}=\{\boldsymbol{x}^{1}_{t},\ldots,\boldsymbol{x}^{N}_{t}\}$ in the multi-batch setting with a batch size of $N$. $\boldsymbol{x}_{c}$ and $\boldsymbol{x}_{t}$ are then sent to the context encoder and target encoder, respectively.
Context Encoder & Target encoder
@@
Learnable Empty Token and Mask Token
@@
where $P_{n}$ is the subset containing all indices of masked empty BEV grids, and $Q_{n}$ is the subset containing all indices of masked non-empty BEV grids in the $n$-th input in a multi-batch setting. We also denote $K_{n}$ as the subset containing all indices of masked non-empty BEV grids in the $n$-th input.
In order to avoid representation collapse, we apply an additional variance regularization loss proposed by (Bardes, Ponce, and LeCun 2022):
@@
where the function $v(\cdot)$, which uses an arbitrary input 2D embedding matrix $Y\in\mathbb{R}^{M\times C}$ recording an arbitrary number $M$ of embeddings with a fixed embedding dimension $C$, is defined in (Bardes, Ponce, and LeCun 2022):
@@
The parameters of the context encoder, the predictor, as well as the learnable mask token and the learnable empty token, are all updated by gradient descent using Figure 4. Meanwhile, the parameters of the target encoder are updated through a moving average of the context encoder’s parameters, $\phi\leftarrow\eta\phi+(1-\eta)\theta$, to further avoid representation collapse.
Experiments
To evaluate our pre-training method, We utilize the OpenPCDet framework (Team 2020) to pre-process the datasets. We conduct experiments on three datasets of increasing scale: the small‑scale KITTI3D (Geiger, Lenz, and Urtasun 2012) (7k frames), the medium‑scale Waymo (Sun et al. 2020) (30k frames for the 20% subset and 150k frames for the full set), and the large‑scale ONCE (Mao et al. 2021) (100k, 500k, and 1M frames for the small, medium, and large splits). We then fine‑tune on a LiDAR 3D object detection task using different network architectures, including SECOND (Yan, Mao, and Li 2018), PV‑RCNN (Shi et al. 2020), and CenterPoint (Yin, Zhou, and Krahenbuhl 2021).
@@
We compare models trained from scratch and reproduce the most recent state-of-the-art (SOTA) self-supervised pre-training methods, Occupancy-MAE (Min et al. 2023) and ALSO (Boulch et al. 2023) that are open-sourced and contain no running issues. Specifically, we reproduce Occupancy-MAE using their default settings to obtain the pre-trained weights and fine-tune the downstream object detector on the KITTI3D, Waymo, and ONCE 100k datasets. For ALSO, we directly fine-tune the released pre-trained encoder weights for the downstream object detector on the KITTI3D dataset. But for ALSO’s ONCE‑100k experiments, we observe that the released pre‑trained weights lead to significant negative transfer in three random runs, therefore we compare our results with those reported in their paper. The comparison methods use the same fine-tuning settings as ours. For all fine-tuning experiments, we fine-tune the model through three independent runs in default settings and report the best performance, in the same manner as (Sautier et al. 2024) reports their results. The reproduced results may vary from the original papers. Additionally, for the largest ONCE dataset, we include comparisons against the popular general self-supervised learning algorithms BYOL (Grill et al. 2020), PointContrast (Xie et al. 2020), SwAV (Caron et al. 2020), DeepCluster (Caron et al. 2018), and DepthContrast (Zhang et al. 2021), using results from ONCE’s official benchmark website. We also include a comparison to ProposalContrast (Yin et al. 2022), using results reported in (Min et al. 2023).
Due to space constraints, additional details, such as dataset descriptions, hyperparameter settings, training protocols, visualizations, further ablation studies on masking ratios, predictor architecture, learnable empty/mask tokens, and variance regularization, as well as detailed comparisons of pre-training efficiency with other methods, etc., can be found in the supplementary material.
Pre-training Efficiency
Unlike Occupancy‑MAE, which uses computationally expensive dense 3D convolutions to reconstruct invisible regions, AD‑L‑JEPA employs a joint‑embedding predictive architecture at the BEV level and omits those layers, resulting in $2.8\times$$3.4\times$ lower GPU memory usage and $2.7\times$ fewer GPU hours for pre‑training on the 20% and 100% splits of the Waymo dataset, and $3.1\times$$4\times$ lower GPU memory usage and $1.9\times$ fewer GPU hours for pre‑training on the ONCE 100k split.
Downstream Fine-tuning Performance
We first conduct self-supervised pre-training on KITTI3D without any labels and then fine-tune the network on the labeled training data from the same KITTI3D dataset using the widely used SECOND (Yan, Mao, and Li 2018) and PV-RCNN (Shi et al. 2020) methods. As shown quantitatively in Table 1, our self-supervised pre-training approach significantly outperforms the baseline that corresponds to training from scratch and also surpasses other competitor algorithms.
@@
Table 1: 3D detection results on the KITTI3D validation set, reported with the AP (%) metric. The models are pre-trained with and fine-tuned on the KITTI3D dataset.
@@
Centerpoint - AP metric
No pre-training 63.28 63.95 66.77 64.67
Occupancy-MAE, 20% 63.20 64.20 67.20 64.87 +0.20
Occupancy-MAE, 100% 63.53 64.73 67.77 65.34 +0.67
AD-L-JEPA (ours), 20% 63.18 64.35 67.68 65.07 +0.40
AD-L-JEPA (ours), 100% 63.58 64.58 68.07 65.41 +0.74
Centerpoint - APH metric